OpenStudy (anonymous):

Find the unit rate 70mi/5h

OpenStudy (anonymous):

The anwser is 14mi/1h SHOW HO TO GET THIS ANWSER

OpenStudy (agreene):

70/5 = 14 mi/h or 14mph if you prefer that notation.

jimthompson5910 (jim_thompson5910):

70 mi/5 h Divide each component by the GCF 5 to get 70/5 = 14 5/5 = 1 So 70 mi/5 h = 14 mi/1 hr

OpenStudy (anonymous):

thanks :) U rock

jimthompson5910 (jim_thompson5910):

np

hero (hero):

70/5 = 14 That's all you need to do